OpenStudy (anonymous):

Evaluate the expression below when g = -6. -3g-2-(-4g+10)=

OpenStudy (anonymous):

write -6 instead of ever g in the expression

OpenStudy (anonymous):

-3*(-6)-2..........

OpenStudy (anonymous):

16?

OpenStudy (anonymous):

the first half is 16, but you have to do complete expression... don't forget about -(-4g+10)

OpenStudy (anonymous):

16*10g?

OpenStudy (alfie):

You just have to plug-in the number (-6) as sery here said. -3g-2-(-4g+10)= with g = -6... -3(-6)-2-(-4(-6)+10) = now you just have to recall that two negative numbers in a product make a positive number. Then you are pretty done. :)

OpenStudy (anonymous):

-18 will be the answer........... alfine